Simpson Company applies overhead on the basis of 200% of direct labor cost. Job No. 305 is charged with $150,000 of direct materials costs and $200,000 of manufacturing overhead. The total manufacturing costs for Job No. 305 is 
1. $500,000 
2. $550,000 
3. $450,000 
4. $350,000
Answer #1
Concepts and reason

Cost of Goods Manufactured Schedule: The cost of goods manufactured is calculated in a statement. For the calculation of cost of goods manufactured it takes the beginning and ending work in process, direct materials, direct labor, and manufacturing overhead into consideration.

Job order costing: It is a method of cost accounting, in which cost is collected and Job order costing. It is a method of cost accounting, in which cost is collected and accumulated for each job, work order, or project separately. Especially the job order costing is followed in organizations where customized goods are produced. The costs of goods manufactured incurred in each job, or project, work order are determined in the Cost of Good Manufactured Schedule.

Direct materials: Direct materials are the raw materials that are directly associated with the production of the goods in the factory. Direct material per unit is the cost of direct material incurred to produce one unit of product.

Direct labor cost: the labor cost includes the wages or salaries paid to employees who physically produce goods or services. Labor refers to the actual work that the employees do to produce products. Labor costs refer to the amount of money that the employer pays the employees to complete the work.

Manufacturing overhead costs: The costs which do not relate directly with the manufacturing of products are referred to as manufacturing overhead costs or indirect costs. Example Insurance cost related to factory operations, Depreciation incurred on factory equipment rent, electricity, wages to workers, and depreciation on manufacturing equipment.

Fundamentals

Inventory: Inventory is the total of three components, raw material, work in process and finished goods and recorded as current asset on the balance sheet. Ending inventory includes the ending balances of the aforesaid components.

Raw materials: Raw materials inventory is the source ingredient of production used to manufacture the finished product.

Work-in-process (WIP): Work-in-process inventory is the partially completed goods, which have to go through other processes before being transformed into finished goods. Manufacturing costs like direct labour costs and direct overhead costs are added to work-in-progress value because these are also the manufacturing or production costs.

Finished goods: Finished goods inventory are the finished or completed goods, which are ready for sale.

Cost of goods manufactured: It is an expense to an organization. Cost of goods manufactured is the cost incurred to manufacture the goods. Cost of goods manufactured will be used to calculate the cost of goods sold. It includes the direct material, direct labor cost and manufacturing overhead.

Cost of goods sold: It is an expense to an organization. Cost of goods sold is the cost incurred to manufacture the goods sold. Cost of goods sold is the cost of beginning inventory, plus cost of additional purchases, minus cost of ending inventory. Therefore, cost of goods sold could be defined as the amount of cost incurred for the products sold.

The formula for total manufacturing cost is shown below:

Totalmanufacturingcost=(Directmaterialused+Directlabor+Manufacturingoverheadapplied){\rm{Total manufacturing cost}} = \left( \begin{array}{l}\\{\rm{Direct material used}}\\\\ + {\rm{Direct labor}} + {\rm{Manufacturing }}\\\\{\rm{overhead applied}}\\\end{array} \right)

Compute direct labor cost:

Directlaborcost=[(Appliedmanufacturingoverheadcost200%)×100]=[($200,000200%)×100]=$100,000\begin{array}{c}\\{\rm{Direct labor cost}}\,{\rm{ = }}\left[ {\left( {\frac{\begin{array}{l}\\{\rm{Applied manufacturing}}\\\\{\rm{ overhead cost }}\\\end{array}}{{{\rm{200\% }}}}} \right) \times 100} \right]\\\\ = \left[ {\left( {\frac{{{\rm{\$ 200,000 }}}}{{{\rm{200\% }}}}} \right) \times 100} \right]\\\\ = \$ 100,000\\\end{array}

Calculate total manufacturing cost for Job No. 305:

TotalmanufacturingcostforJobNo.305}=(Directlaborcost+Directmaterialcost+Manufacturingoverheadapplied)=($100,000+$150,000+$200,000)=$450,000\begin{array}{c}\\\left. \begin{array}{l}\\{\rm{Total manufacturing cost }}\\\\{\rm{for Job No}}.{\rm{305}}\\\end{array} \right\} = \left( \begin{array}{l}\\{\rm{Direct labor cost + }}\\\\{\rm{Direct material cost }}\\\\ + \,{\rm{Manufacturing }}\\\\{\rm{overhead applied}}\\\end{array} \right)\\\\ = \left( \begin{array}{l}\\\$ 100,000 + \$ 150,000\\\\ + \$ 200,000\\\end{array} \right)\\\\ = \$ 450,000\\\end{array}

Ans:

The total manufacturing cost for Job No. 305 is $450,000.